Now suppose the weekly supply function for Kelewele at UPSA campus is QSx = -260 + 10Px – 2Pi




Where QSx is the quantity supplied of Kelewele and Pi is the price of inputs used in preparing Kelewele




f)What is the supply function if the input prices are ¢ 20Gh




g) Graph the supply curve from f).




Compute the equilibrium price and quantity of Kelewele. Suppose authorities at UPSA are concerned that Kelewele sellers at UPSA are exploiting students by charging exorbitant price for their Kelewele so they decree that no one should sell Kelewele above GHȼ 40 per lb


h) What type of price control measure is this?


I) Following the decree, will there be excess demand or excess supply of Kelewele at UPSA? Calculate the excess demand or excess supply?


j) To ensure that price control measure stands, the authorities have to produce the excess demand or buy the excess supply in i) off the market. How much will this cost the authorities?


1
Expert's answer
2021-05-28T09:44:25-0400

f) "Qs_x=-260+P_x-2(20)=260+10P_x-40=-300+10P_x"

g)



h) Equilibrium occurs at demand = supply

i) "900-10P_x=-300+10P_x"

"1200=20P_x"

"60=P_x"

j) "Q_x=900-10(60)=300"